Assume f element of R(alpha) on [a,b] where alpha is of bounded variation on [a,b] .let V(x) denote the total variation of alpha on [a,x] if a less than x less than or equal to b and let V(a) =0 .show that | integral a to b fdalpha| less than or equal to integral a to b |f| dv less than or equal to MV (b).


1
Expert's answer
2021-03-16T07:14:26-0400

Consider,

"|\int_a^bfd(\alpha)|\le \int_a^b|fd(\alpha)|\\
\le\int_a^b|f||d(\alpha)|\\
\le sup\int_a^b |f||d(\alpha)|\\
=\int_a^b |f|sup |d(\alpha)|\\
=\int_a^b|f|dv ; v= sup|\alpha| \\
\le sup \int_a^b|f| dv\\
=\int_a^bsup|f|dv\\
 =sup|f| \int_a^bdv \\
=MV|_a^b ; M= sup |f|[a, b]\\
=M (V(b)-V(a))\\
=M(V(b)-0);since, v(a) =0\\
=MV(b)\\
Therefore, \\
|\int_a^bfd(\alpha)| \le \int_a^b|f|dv\le MV(b)"